let f(g) be defined as f(g)=g^2-g for all values of g. If f(x) =f(x-3), what is rhe value of x
et fle) be defined as f(g) = g? -g for all values of g. If f(x) = f(x-3), what is the value of x? a) 1 b) c) 2 d) 2.5
Answer #1

9) ng- Giren that f(g)- g²-g (4- Fx=-3 → xr-x = (x-1)*(x-7 zx; -6x+1-X3 xx-x+7x : 9+3 6X = 12. 个个介 企个个 → X 2- Hence the value